Find the quotient and remainder of f(x) = x4 -2 divided by p(x) = x-1
Possible answers:
x3 + x2 +1; -1
x3 + x2 + x +1; -1
x3 + x + 1; -1
x3 - x2 - x - 1; -1

Answer by pradhyumna.kumar@gmail.com(62)

Find the quotient and remainder of f(x) = x4 -2 divided by p(x) = x-1
Possible answers:
x3 + x2 +1; -1
x3 + x2 + x +1; -1
x3 + x + 1; -1
x3 - x2 - x - 1; -1

solution is :- x^3 + x^2 + x + 1 ; -1